OpenStudy (anonymous):

"Find the real number k for which 1 + ki, (i = √-1), is a zero of the polynomial z² + kz + 5" I get k=3 or k=-2, but according to the answer sheet k=-2 is the only answer. Can someone explain why?

OpenStudy (anonymous):

Yes, and I get k= 3 Or k=-2

OpenStudy (anonymous):

it says k is real

OpenStudy (anonymous):

However only k=-2 is supposed to be correct

OpenStudy (anonymous):

z² + kz + 5 roots : 1 + ki, and 1 - ki

OpenStudy (anonymous):

sum of roots : 2 = -k

OpenStudy (anonymous):

product of roots : 1 + k^2 = 5 k^2 = 4 k = +2, - 2

OpenStudy (anonymous):

so i am getting k = 2 and -2

OpenStudy (anonymous):

Wouldn't (1+ki) + (1-ki) be 2?

OpenStudy (anonymous):

stgreen can you show your working out? Because I really don't know what to do here :P

OpenStudy (whpalmer4):

\[(1+ki)^2+k(1+ki)+5=0\] \[(1+2ki-k^2)+k+k^2i+5=0\] To get rid of terms in \(i\) \[2ki=-k^2\]

OpenStudy (whpalmer4):

Uh should be \[2ki=-k^2i\]

OpenStudy (whpalmer4):

Then \(k=-2\)

OpenStudy (anonymous):

I get everything except how you made the jump from (1+2ki−k2)+k+k2i+5=0 to 2ki=−k2i. Can you explain that part?

OpenStudy (anonymous):

Aaa right I get it now, thanks a bunch!

OpenStudy (whpalmer4):

Great! Glad you were able to figure it out...it's a handy trick! Of course you have to verify that everything else sums to 0, but if that didn't, you wouldn't have a solution.